# Polarizable Water Model with Ab Initio Neural Network Dynamic Charges and Spontaneous Charge Transfer

### v1

@@ Abstract
Simulating water accurately has been a challenge due to the complexity of describing polarization and intermolecular charge transfers. Quantum mechanical (QM) electronic structures provide an accurate description of polarization in response to local environments, which is nevertheless too expensive for large water systems. In this study, we have developed a polarizable water model integrating Charge Model 5 atomic charges at the level of second-order Møller-Plesset perturbation theory, predicted by an accurate and transferable Charge Neural Network (ChargeNN) model. The spontaneous intermolecular charge transfer has been explicitly accounted for, enabling a precise treatment of hydrogen bonds and out-of-plane polarization. Our ChargeNN water model successfully reproduces various properties of water in gas, liquid and solid phases. For example, ChargeNN correctly captures the hydrogen-bond stretching peak and bending-libration combination band, which are absent in the spectra using fixed charges, highlighting the significance of accurate polarization and charge transfers. Finally, the molecular dynamical simulations using ChargeNN for liquid water and a large water droplet with a $\sim$4.5 nm radius reveal that the strong interfacial electric fields are concurrently induced by the partial collapse of the hydrogen-bond network and surface-to-interior charge transfers. Our study paves the way for QM-informed force fields, aiming for large-scale molecular simulations with high accuracy.
@@ 1 Introduction
Water is the most common solvent for chemical reactions, and also plays a prominent role in understanding the functions of nano materials1 and biomolecules2, 3. Despite extensive study, the mechanisms behind some of its anomalous properties remain elusive. The polarization effect, which is the electron rearrangement in response to different local environments, has been shown to profoundly influence both static and dynamic properties of water.4, 5, 6, 7 For examples, the induced polarization by electric field is considered to be the major source of the nonadditive cooperativity in water hydrogen bonding interactions.8, 9 Moreover, the considerable intermolecular charge transfers through the water hydrogen bonds make the hydrogen bonds much stronger than the normal noncovalent bonds and possess ”covalency”.10 In recent years, microdroplets have been shown to accelerate chemical reactions by up to 10 million times compared to bulk solutions, widely attributed to the strong electric fields formed on their surfaces. One promising origin of the interfacial electric fields is the large number of H2Oδ+/H2Oδ- radical pairs on the microdroplet surface created by the intermolecular charge transfers11, 12, 13, 14. A recent second-order Møller-Plesset perturbation theory (MP2) based molecular dynamics (MD) study indicates that substantial and inhomogeneous interfacial water charge transfers dramatically promote the reactions of Criegee intermediates with water. 15
@@
Atomic partial charge is a qualitative representation of the environment-dependent electron distribution for understanding polarization effects and rationalizing chemical phenomena.16, 17, 18, 19, 20. Nevertheless, partial charge is neither experimentally nor theoretically observable. In quantum chemistry, the atomic charges vary significantly with different electron density partition methods, such as Mulliken21, Löwdin22, natural bond orbital23, atoms in molecules24, Hirshfeld25, etc. Notably, the Hirshfeld population analysis precisely describes the density redistribution during bonding.25 Furthermore, the charge model 5 (CM5) corrects the underestimated Hirshfeld charges to achieve accurate dipole moments26, which has been shown to provide accurate electronic distribution and to be insensitive to the choice of basis sets and theories.27 Calculating partial charges is computationally demanding using density functional theory (DFT) or higher level correlated wavefunction methods (e.g. MP2 and coupled cluster singles and doubles (CCSD)), making it cost-prohibitive for large molecular systems. In recent years, efforts have been made to map the atomistic local environment descriptors onto QM atomic charges for accurate and efficient prediction of atomic charges by harnessing cutting-edged machine learning (ML) technologies.28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47 However, these charge models were mainly tested on single molecules and small clusters, casting doubt on their accuracy when applied to larger molecular systems. A recent study reports a kernel ridge regression (KRR) model accurately predicts iterative Hirshfeld atomic charges48 of water molecules in bulk liquid represented by QM/molecular mechanical (QM/MM) water clusters.32 Remarkably, the hydrogen-bond stretch infrared (IR) peak is successfully retrieved using dynamic ML charges, whereas this weak peak disappears for fixed charges, marking the importance of polarization and intermolecular charge transfers in simulating liquid water. 32 However, this model has several limitations. Firstly, the training and testing were conducted with an identical system (23 QM water molecules with 1977 MM water molecules), leaving the model’s generalizability to larger QM regions uncertain. Secondly, the charge model was specifically tailored for water in bulk solution, which makes it inadequate for predicting charges for interfacial water molecules that experience more complex local environments. Finally, while ML charges were used to compute dipole moments and IR spectra using trajectories prepared with TIP4P/2005, they have not been integrated into a water model for simulating polarization in molecular dynamics.
@@
Water has been challenging molecular simulations for decades.49 Since the introduction of the rigid non-polarizable water models with fixed charges50, 51, 52, they have been substantially improved by refining parameters 53, 54, 55, 56, 57, introducing intramolecular flexibility58, 59, 60, and integrating implicit polarizability61. Moreover, tremendous efforts have been made to recover the explicit polarization effects. Some polarizable models incorporate induced multipole in response to the local electric fields using atomic polarizability tensors 62, 63, 64, 65, 66, 67, 68, 69, 70 or drude oscillators71, 72, 73, 74, 75. Unlike the induced multipole models which still use fixed charges, the charge-flow models enables geometry-dependent charge redistribution76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91. Although charge-flow polarizable models appear enticing due to their explicit incorporation of charge exchange between atoms, they suffer from several deficiencies, including high computational costs associated with iteratively updating charges, non-linear computational scaling of the polarizability with system sizes, inadequate description of intermolecular charge transfer, and a lack of out-of-plane polarization.92, 93 Furthermore, despite extensive fine-tuning of the charge functional parameters to fit the DFT electrostatic potentials, the absence of spontaneous quantum mechanical information during the solution of dynamical charges may result in questionable electron density arrangements. Beyond the empirical force fields, energy potentials from first-principles for large water systems can be built by employing many-body expansion (MBE) for potentials of small water fragments obtained through fitting with functionals or machine learning.94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110 Nevertheless, constructing an accurate potential requires a tremendous amount of training data to cover the vast chemical space. The applicability of potentials is severely limited due to their system-specific nature that does not generalize well to unseen molecules. Furthermore, important electronic properties such as electron density and multipole moments are absent in the potentials.
@@
In this study, we present a novel polarizable water model that incorporates dynamical atomic charges and charge transfers at the ab initio MP2 level of theory. The charges are produced using a deep neural network charge model called ChargeNN that is equipped with the interaction classified functions (ICFs) to faithfully depict the local environments. Trained on diverse (H2O)25 structures sampled from MD trajectories, the ChargeNN is capable of accurately predicting charges for much larger water clusters, for example, (H2O)190, as compared to computed MP2 charges. The ChargeNN has also been integrated into a water model for dynamical polarization simulations. The intermolecular charge transfer has been explicitly accounted for using machine-learned pair-wise water charge transfers, which effectively captures the out-of-plane polarization. We have implemented analytical energy gradients in the periodic boundary condition (PBC) within the shared-memory message passing interface (MPI) framework for efficient MD simulations. Our results validate the capability of ChargeNN-derived water model to reproduce a variety of water properties in excellent agreement with experimental data, for instance, the model generalizes well to a few thousand water molecules in unit cell for predicting the liquid water properties. Finally, the statistical analysis of liquid water and large water microdroplets using ChargeNN reveals that the partial breakage of hydrogen-bond networks, along with water charge transfers, concurrently promotes the layer electric fields at the air/water interface. The ChargeNN water model is thus of broad applicability in studying condensed ice, liquid water, nano-phase, large microdroplets, and solvent interactions with materials and biomolecules.
@@ 2.1 Deep Neural Networks for Atomic Charges
Tremendous efforts have been made to design descriptors of local atomic environment for successful applications.105, 111, 106, 107, 108 However, the existing descriptors omit the types of atomic interactions that may have distinct weights to the total local interactions. To incorporate such important information, we propose an ICF to describe the interaction-specific ($\alpha\beta$) contributions to the charge of atom $i$ of the atom type $\alpha$, which satisfies translational, rotational and mutational invariance:
$\begin{split}f^{\alpha\beta}_{i}={}&\sum_{j\in\beta}e^{-\frac{r_{ij}^{2}}{k_{f}}}s_{c}(r_{ij}),\textrm{ where }i\in\alpha,\\ s_{c}(r_{ij})={}&\begin{cases}0.5\left[\text{cos}\left(\frac{\pi r_{ij}}{r_{c}}\right)+1\right],&r_{ij}\leq r_{c}\\ 0,&r_{ij}>r_{c}\end{cases}\end{split}$ (1)
In the above formula, the parameter $k_{f}$ controls the interaction decay with respect to the atomic distance. $s_{c}$ denotes the cutoff function that enforces a smooth transition of atoms $j$ entering and exiting the selection region for the atom $i$. We tested the charge convergence for both bulk water and air/water interface, and found that 24 surrounding water molecules sufficiently form the local environment to produce charges comparable to those surrounded by 127 molecules, as demonstrated in Figure S1. The selected cutoff $r_{c}$ was accordingly set to be 4.4 Å, which is the minimum radius of multiple (H2O)25 droplets extracted from well-equilibrated liquid water boxes.
@@
$\mathbf{F}_{iI}=\{f_{iI}^{\alpha_{I}\beta_{I}},f_{iI}^{\alpha_{I}\beta_{J}},f_{jI}^{\alpha_{I}\beta_{J}}\}$ (2)
where the intra- ($f_{iI}^{\alpha_{I}\beta_{I}}$) and intermolecular ICFs ($f_{iI}^{\alpha_{I}\beta_{J}}$) are separately placed for detailing interaction types. In addition, the intermolecular ICFs of other atoms in the same molecule ($f_{jI}^{\alpha_{I}\beta_{J}}$) are also appended to the feature set to account for the the perturbation arising from the interactions for the neighboring atoms, which notably improves the prediction accuracy. Three decay parameters $k_{f}$ were employed for each ICF in eq. 2 to simulate varying interactions with the atomic distance. To better handle the distinct charge ranges and distributions, two separate neural networks were specifically trained for the charges of oxygen and hydrogen.
@@
$\begin{split}\mathbf{F}_{IJ}={}&\{f^{\alpha_{I}\beta_{J}}\},\\ f^{\alpha_{I}\beta_{J}}={}&\sum_{\begin{subarray}{c}i\in\alpha I\\ j\in\beta J\end{subarray}}e^{-\frac{r_{ij}^{2}}{k_{f}}}s^{\mathrm{CT}}_{c}(r_{ij}).\\ \end{split}$ (3)
@@
$A(x)=\frac{x}{1+e^{-\beta x}}$ (4)
The Swish function is differentiable and ensures smooth charge functions for new molecular configuration, which avoids discontinuity on the potential energy surface. Following multiple trials, we discovered that utilizing $\beta=3$ allows the activation function to retain good linearity around $x=0$, enhancing the learning performance while circumventing erratic charge functions caused by overfitting.
The CM5 atomic charges of (H2O)25 were utilized to train the neural networks. Reference charges were computed using the low-scaling third-order MBE orbital-specific virtual MP2 (MBE(3)-OSV-MP2) method112 with the cc-pVTZ basis set, as MP2 has been shown to accurately predict the energetic and electronic properties of water113, 114, 115. To benchmark the charge model, charges were prepared for 20000 (H2O)25, 200 (H2O)32, 50 (H2O)64, 10 (H2O)64 and 1 (H2O)190 cluster conformations that were randomly sampled from the MD trajectories using SWM4-NDP with harmonic bonds and angles on OPENMM116, 117.

### v3

@@ Abstract
Simulating water accurately has been a challenge due to the complexity of describing polarization and intermolecular charge transfer. Quantum mechanical (QM) electronic structures provide an accurate description of polarization in response to local environments, which is nevertheless too expensive for large water systems. In this study, we have developed a polarizable water model integrating Charge Model 5 atomic charges at the level of second-order Møller-Plesset perturbation theory, predicted by an accurate and transferable Charge Neural Network (ChargeNN) model. The spontaneous intermolecular charge transfer has been explicitly accounted for, enabling a precise treatment of hydrogen bonds and out-of-plane polarization. Our ChargeNN water model successfully reproduces various properties of water in gas, liquid and solid phases. For example, ChargeNN correctly captures the hydrogen-bond stretching peak and bending-libration combination band, which are absent in the spectra using fixed charges, highlighting the significance of accurate polarization and charge transfer. Finally, the molecular dynamical simulations using ChargeNN for liquid water and a large water droplet with a $\sim$4.5 nm radius reveal that the strong interfacial electric fields are concurrently induced by the partial collapse of the hydrogen-bond network and surface-to-interior charge transfer. Our study paves the way for QM-polarizable force fields, aiming for large-scale molecular simulations with high accuracy.
@@ 1 Introduction
As an active reagent, water plays a crucial role in regulating biomolecular functions, for instance, to stabilize their structures and facilitate dynamic interactions in proteins and nucleic acids.1, 2 Water also exhibits anomalous characteristics for nanomaterials, such as unique structural arrangements and fast diffusion rates within confined nanotubes compared to the bulk phase3. The polarization effect, which is the electron rearrangement in response to different local environments, has been shown to profoundly influence both static and dynamic properties of water.4, 5, 6, 7 For example, the induced polarization by electric field is considered to be the major source of the nonadditive cooperativity in water hydrogen bonding interactions.8, 9 Moreover, the considerable electron transfer between water molecules enhances the proton donor-acceptor orbital interactions and hence the covalency, directionality as well as relevant spectroscopic properties of water hydrogen bonds.10 In recent years, microdroplets have been shown to accelerate chemical reactions by up to 10 million times compared to bulk solutions, widely attributed to the strong electric fields formed on their surfaces. One promising origin of the interfacial electric fields is the large number of H2Oδ+/H2Oδ- radical pairs on the microdroplet surface created by the intermolecular charge transfer11, 12, 13, 14. A recent second-order Møller-Plesset perturbation theory (MP2) based molecular dynamics (MD) study indicates that substantial and inhomogeneous interfacial water charge transfer dramatically promotes the reactions of Criegee intermediates with water. 15
@@
Atomic partial charge is a qualitative representation of the environment-dependent electron distribution for understanding polarization effects and rationalizing chemical phenomena.16, 17, 18, 19, 20. Nevertheless, partial charge is neither experimentally nor theoretically observable. In quantum chemistry, atomic charges vary significantly with different electron density partition methods, such as Mulliken21, Löwdin22, natural population analysis (NPA)23, quantum theory of atoms in molecules (QTAIM)24, Hirshfeld25, etc. Notably, the Hirshfeld population analysis precisely describes the density redistribution during bonding.25 Furthermore, the charge model 5 (CM5) corrects the underestimated Hirshfeld charges to achieve accurate dipole moments26, which has been shown to provide accurate electronic distribution and to be insensitive to the choice of basis sets and theories.27 Calculating accurate electron densities to produce partial charges is computationally demanding using density functional theory (DFT) or higher level correlated wavefunction methods (e.g. MP2 and coupled cluster singles and doubles (CCSD)), making it cost-prohibitive for large molecular systems. In the past decades, efforts have been made to develop machine learning (ML) charges by mapping the atomistic local environment descriptors onto QM reference atomic charges.28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50 However, these charge models were mainly tested on single molecules and small clusters, casting doubt on their accuracy when applied to larger molecular systems. A recent study reports a kernel ridge regression (KRR) model accurately predicts iterative Hirshfeld atomic charges51 of water molecules in bulk liquid represented by QM/molecular mechanical (QM/MM) water clusters.35 Remarkably, the hydrogen-bond stretch infrared (IR) peak is successfully retrieved using dynamic ML charges, whereas this weak peak disappears for fixed charges, marking the importance of polarization and intermolecular charge transfer in simulating liquid water. 35 However, this model has several limitations. Firstly, the training and testing were conducted with an identical system (23 QM water molecules with 1977 MM water molecules), leaving the model’s generalizability to larger QM regions uncertain. Secondly, the charge model was specifically tailored for water in bulk solution, which makes it inadequate for predicting charges for interfacial water molecules that experience more complex local environments. Finally, while ML charges were used to compute dipole moments and IR spectra using trajectories prepared with TIP4P/2005, they have not been integrated into a water model for simulating polarization in molecular dynamics.
@@
Simulating water has been challenging for decades.52 Since the introduction of the rigid non-polarizable water models with fixed charges53, 54, 55, they have been substantially improved by refining parameters 56, 57, 58, 59, 60, introducing intramolecular flexibility61, 62, 63, and integrating implicit polarizability64. Moreover, tremendous efforts have been made to recover the explicit polarization effects. Some polarizable models incorporate induced multipoles in response to the local electric fields using atomic polarizability tensors 65, 66, 67, 68, 69, 70, 71, 72, 73, 28, 29, 74, 75, 76, 77 or Drude oscillators78, 79, 80, 81, 82. In contrast, the charge-flow models enable geometry-dependent charge redistribution83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98. Although charge-flow polarizable models appear enticing due to their explicit incorporation of charge exchange between atoms, they suffer from several deficiencies, including high computational costs associated with iteratively updating charges, non-linear computational scaling of the polarizability with system sizes, inadequate description of intermolecular charge transfer, and a lack of out-of-plane polarization.99, 100 Furthermore, despite extensive fine-tuning of the charge functional parameters to fit the DFT electrostatic potentials, the absence of spontaneous quantum mechanical information during the solution of dynamical charges may result in questionable electron density arrangements. Beyond the empirical force fields, energy potentials from first-principles for large water systems can be built by employing many-body expansion (MBE) for potentials of small water fragments obtained through fitting with functionals or machine learning.101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117 Nevertheless, constructing an accurate potential requires a tremendous amount of training data to cover the vast chemical space. The applicability of potentials is severely limited due to their system-specific nature that does not generalize well to unseen molecules. Furthermore, important electronic properties such as electron density and multipole moments are absent in the potentials. Notably, the QM-polarizable force fields that incorporate instantaneous QM-level polarizability enabled by machine-learning, such as FFLUX using machine learned QTAIM atomic multipoles, have been successfully applied to various chemical systems.28, 29, 118, 74, 75, 76, 77
@@
In this study, we present a novel QM-polarizable water model that incorporates dynamical atomic charges and charge transfer exclusively trained on QM data acquired at the ab initio MP2 level of theory. The charges are produced using a deep neural network charge model called ChargeNN. A set of Interaction Classified Functions (ICFs) was employed as the ChargeNN features, explicitly capturing the types of atomic interactions to accurately depict local environments. Trained on diverse (H2O)25 structures sampled from MD trajectories, the ChargeNN is capable of accurately predicting charges for much larger water clusters, for example, (H2O)190, as compared to computed MP2 charges. The ChargeNN has also been integrated into a water model for dynamical polarization simulations. The intermolecular charge transfer has been explicitly accounted for using machine-learned pair-wise water charge transfer, which effectively captures the out-of-plane polarization. We have implemented analytical energy gradients in the periodic boundary condition (PBC) within the shared-memory message passing interface (MPI) framework for efficient MD simulations. Our results validate the capability of ChargeNN-derived water model to reproduce a variety of water properties in excellent agreement with experimental data, for instance, the model generalizes well to a few thousand water molecules in unit cell for predicting the liquid water properties. Finally, the statistical analysis of liquid water and large water microdroplets using ChargeNN reveals that the partial breakage of hydrogen-bond networks, along with water charge transfer, concurrently promotes the layer electric fields at the air/water interface. The ChargeNN water model is thus of broad applicability in studying condensed ice, liquid water, nano-phase, large microdroplets, and solvent interactions with materials and biomolecules.
@@ 2 Methods
2.1 Charge Model 5
Charge Model 5 (CM5)26 based on Hirshfeld population analysis (HPA)25 was implemented for third-order many-body-expansion orbital-specific-virtual MP2 (MBE(3)-OSV-MP2) to investigate the water polarization and charge transfer in response to the variation in the local environment. The Hirshfeld atomic charges ($q^{\mathrm{HPA}}_{i}$) are obtained by integrating the electron density within the region surrounding each atom,
$q^{\mathrm{HPA}}_{i}=Z_{i}-\int\mathrm{d}\mathbf{r}\frac{\rho^{0}_{i}(\mathbf{r})}{\sum_{j}\rho^{0}_{j}(\mathbf{r})}\rho(\mathbf{r}),$ (1)
where $Z_{i}$ is the nuclear charge. $\rho^{0}(\mathbf{r})$ and $\rho(\mathbf{r})$ denote the electron densities of the promolecule and real molecule at position $\mathbf{r}$, respectively. In contrast to the hard-boundary partitioning approaches like QTAIM119, 120, where electron density at each grid point is attributed to a single atom, the boundary-less Hirshfeld population analysis allows electron density to contribute to multiple atoms. It provides a clear partitioning of the electron density and accurately describes the density redistribution from promolecule to real molecule during bonding. Nevertheless, Hirshfeld charges are significantly underestimated due to the Hirshfeld weighting factor, which causes the atomic population to closely resemble that of the isolated atom. 51
By introducing a unified set of parameters optimized by fitting to the experimental or DFT dipole moments of 614 molecules, the CM5 model maps the underestimated Hirshfeld charges onto a new set of charges, providing a more accurate description of the electrostatic potential:
$\begin{split}q^{\mathrm{CM5}}_{i}={}&q^{\mathrm{HPA}}_{i}+\sum_{j\neq i}A_{ij}B_{ij},\\ B_{ij}={}&\text{exp}\left[-\alpha(r_{ij}-R_{Z_{i}}-R_{Z_{j}})\right],\end{split}$ (2)
where $A_{ij}$ and $\alpha$ are the CM5 model coefficients to be optimized. Constant $R_{Z}$ denotes the atomic covalent radius. CM5 has been shown to deliver precise electronic charge distributions and be insensitive to different theoretical levels and basis sets.26, 27. As demonstrated in Figure 1a, CM5 charges outperform iterative Hirshfeld (Hirshfeld-I), NPA, and QTAIM charges in reproducing the dipole moments of water clusters containing 1-20 molecules. Furthermore, CM5 exhibits numerically consistent dipole moments between cc-pVTZ and aug-cc-pVTZ basis sets, as shown in Figure 1b. The CM5 charges in this work were obtained from the MBE(3)-OSV-MP2/cc-pVTZ relaxed density matrix.
Figure 1: Water dipole moments derived from charges (a) Water dipole moments obtained from CM5, Hirshfeld-I51, NPA121 and QTAIM charges, using electron density of MP2/aug-cc-pVTZ computed with Gaussian 16122. CM5 and NPA charges were obtained with Gaussian 16, while Multiwfn123, 124 was used to compute Hirshfeld-I and QTAIM charges. (b) Water dipole moments obtained from CM5 charges, computed with MBE(3)-OSV-MP2 using aug-cc-pVTZ and cc-pVTZ basis sets. Experimental structure of water monomer and structures from Cambridge water clusters125 were used for this test. (H2O)6c and (H2O)6p refer to cage and prism conformers of water hexamers, respectively. Experimental dipole moment of the water monomer is from ref.126.
@@ 2.2 Deep Neural Networks for Atomic Charges
Tremendous efforts have been made to design descriptors of local atomic environments for building machine learning models of energy potentials and atomic properties.112, 127, 128, 113, 114, 115 For instance, the well known radial ($G_{i}^{\text{rad}}$) and angular ($G_{i}^{\text{ang}}$) atom-centered symmetry functions (ACSFs)127 write
$\begin{split}G_{i}^{\text{rad}}={}&\sum_{j}e^{-\eta(r_{ij}-r_{s})^{2}}s_{c}(r_{ij}),\\ G_{i}^{\text{ang}}={}&2^{1-\zeta}\sum_{\begin{subarray}{c}j,k\neq i\end{subarray}}\left(1+\lambda\cos\theta_{ijk}\right)^{\zeta}\cdot e^{-\eta\left(r_{ij}^{2}+r_{ik}^{2}+r_{jk}^{2}\right)}\\ {}&\cdot s_{c}\left(r_{ij}\right)s_{c}\left(r_{ik}\right)s_{c}\left(r_{jk}\right),\end{split}$ (3)
where $\eta$, $r_{s}$, $\zeta$ and $\lambda$ are tunable parameters. $s_{c}$ denotes the cutoff function that enforces a smooth transition of atoms $j$ entering and exiting the selection region for the atom $i$
$s_{c}(r_{ij})=\begin{cases}0.5\left[\text{cos}\left(\frac{\pi r_{ij}}{r_{c}}\right)+1\right],&r_{ij}\leq r_{c}\\ 0,&r_{ij}>r_{c}\end{cases}.$ (4)
However, ACSF descriptors omit the dependence of the atomic interactions on element type, which can significantly impact training performance and data efficiency.
To incorporate the information of atomic interaction types, we propose an interaction classified function (ICF) to describe the interaction-specific ($\alpha\beta$) contributions to the charge of atom $i$ of the atom type $\alpha$, which satisfies translational, rotational and permutational invariance but requires no angular ACSF parameterization for angularly independent atomic charges:
$\begin{split}f^{\alpha\beta}_{i}={}&\sum_{j\in\beta}e^{-\frac{r_{ij}^{2}}{k_{f}}}s_{c}(r_{ij}),\textrm{ where }i\in\alpha.\\ \end{split}$ (5)
In the above formula, a single parameter $k_{f}$ controls the interaction decay with respect to the atomic distance, effectively alleviating the heavy process of the parameter tuning. Despite their simplicity, the radial ICFs have been found to sufficiently describe local interactions needed for reproducing accurate atomic charges with excellent data efficiency (see section 3.1). To determine the atom selection cutoff ($r_{c}$) in eq. 4, we tested the charge convergence for both bulk water and air/water interface, and found that 24 surrounding water molecules sufficiently form the local environment to produce charges comparable to those surrounded by 127 molecules, as demonstrated in Figure S1. The selected cutoff $r_{c}$ was accordingly set to be 4.4 Å, which is the minimum radius of multiple (H2O)25 droplets extracted from well-equilibrated liquid water boxes.
@@
$\mathbf{F}_{i_{I}}=\{f_{i_{I}}^{\alpha_{I}\beta_{I}},f_{i_{I}}^{\alpha_{I}\beta_{J}},f_{j_{I}}^{\alpha_{I}\beta_{J}}\}.$ (6)
where the intra- ($f_{i_{I}}^{\alpha_{I}\beta_{I}}$) and intermolecular ICFs ($f_{i_{I}}^{\alpha_{I}\beta_{J}}$) are separately placed for detailing interaction types. In addition, the intermolecular ICFs of other atoms in the same molecule ($f_{j_{I}}^{\alpha_{I}\beta_{J}}$) are also appended to the feature set to account for the perturbation arising from the interactions for the neighboring atoms, which notably improves the prediction accuracy. Three decay parameters $k_{f}$ were employed for each ICF in eq. 6 to simulate varying interactions with the atomic distance, as illustrated in Figure S2. To better handle the distinct charge ranges and distributions, two separate neural networks were specifically trained for the charges of oxygen and hydrogen.
@@
$\begin{split}\mathbf{F}_{IJ}={}&\{f^{\alpha_{I}\beta_{J}}\},\\ f^{\alpha_{I}\beta_{J}}={}&\sum_{\begin{subarray}{c}i\in\alpha_{I}\\ j\in\beta_{J}\end{subarray}}e^{-\frac{r_{ij}^{2}}{k_{f}}}s^{\mathrm{CT}}_{c}(r_{ij}).\\ \end{split}$ (7)
@@
The CM5 atomic charges of (H2O)25 were utilized to train the neural networks. Reference charges were computed using the low-scaling MBE(3)-OSV-MP2 method129 with the cc-pVTZ basis set, as MP2 has been shown to accurately predict the energetic and electronic properties of water130, 131, 132. To benchmark the charge model, charges were prepared for 20000 (H2O)25, 200 (H2O)32, 50 (H2O)64, 10 (H2O)64 and 1 (H2O)190 cluster conformations that were randomly sampled from the equilibrated MD trajectories using SWM4-NDP with harmonic bonds and angles on OpenMM133, 134. The $NVT$ trajectories were simulated in the gas phase at 298.15 K, with a time step of 1 femtosecond.
Four hidden layers are employed to effectively capture intricate relationships between features and atomic charges, which also mitigates overfitting issues. The Swish function135 is adopted as the activation function:
$A(x)=\frac{x}{1+e^{-\beta x}}.$ (8)
The Swish function is differentiable and ensures smooth charge functions for new molecular configuration, which avoids discontinuity on the potential energy surface. Following multiple trials, we discovered that utilizing $\beta=3$ allows $A(x)$ to retain good linearity around $x=0$, enhancing the learning performance while circumventing erratic charge functions caused by overfitting. The mean squared error (MSE) was chosen as the loss function. Among the 20,000 configurations of (H2O)25, 8,000 were randomly selected for training, with 50 used as the validation data set. The remaining data points were used for evaluating the prediction performance. After extensive testing, we applied the “early stopping” technique with 200 epochs to ensure satisfactory loss convergence and to prevent overfitting.